flep:

How do you evaluate the integral ∫𝜋0sin100(𝑥)𝑑𝑥 ?

Tranquility:

Do you really need help with this? You've been asking basic arithmetic questions this past hour and now you're asking a calculus question...

Tranquility:

Would you mind using the draw tool below to write out the integral?

Tranquility:

Is this the question \[\int\limits_{0}^{\pi} \sin100x ~dx\]

Tranquility:

Can you clarify the question then?

Tranquility:

You can use the equation tool to also write out the question. You should know \[\int \sin u ~du = -\cos u + C\]